DECEMBER PLAN: Bring up ALk to 8.6 and after that is achievced I will begin to dose 5ml/day B-Ionic solution to keep alkalinity stable and I will monitor Calcium to see when I need to dose. To remove Cyano bacteria I will Keep up a weekly water change schedule and manually remove what I can. Testing will be done bi-weekly before waterchange and once after to monitor progress.

As of Right now:
Alk: 8.23
Ca: 435
View November 29, 2020 18:32
Removed rock with aptasia and did 10gal WC. Detritus was syphoned out and rocks were brushed off.
.
Found Problem at RODI system. System was adding 200 TDS into tank the entire time..... Replaced Both Membranes 11/13/2020. Found DI Resin Cartrages to be installed upside down.. Ordered New DI Resin. Top Off water now has 000 TDS:).
.
Goal is to not have Cyano and Dinos with weekly WC's.
View November 15, 2020 02:45
Replaced RODI Unit Membranes and fixed DI Resin filters. 0 TDS Comming out of the DI Resin.
Need to do 20% WC Tomorrow morning.
View November 14, 2020 06:42
Increased T5 Photoperiod from 10am-7pm to 9am-9pm for an increase of 3 Total hours.
View October 19, 2020 16:03
PC Rainbow browing out but other coral look okay.
1)Nitrates 12, Phosphate 0.08
------>Turned on skimmer, replaced filter sock.
2) possible not enough light??
------>Increased Blue, Royal Blue, and deep blue to 90% and white to 15%
.
.
Red Algae becoming an issue
.......need to manually remove during next WC.
.
.
Aptasia in tank x10
.......Get Aptasia X back from dad.
.
.
Dinos seem to be suppressed with the addition of the UV Sterilizer. :)
View October 18, 2020 22:24
Mounted all frags onto rocks.
Added Another Hydor pump 1000+ gph
View September 27, 2020 18:59
EVS B-Ionic Nitrate works great. 1ml/day for 4 days started 9/16 (2ppm). Tested on the 5th day 9/20 (8ppm) and stopped dosing. Retested on 9/27 (4ppm) Nitrate..... So After Next 20 % Water Change try dosing 1ml/day for 4 days and retest.
.
.
.
Increased AI 52 Light Intensity
White 8-->10
Blue 80-->85
Royal 80-->85
.
.
.
Overall slow growth in tank but polyp extension on all Acropora.
Alkalinity uptake seems steady at about 0.2 dkh.
Red/Purple slime or Coraline on Rocks.
Film Algae Is abundant.
Dinos/Air Bubbles are still present in tank.
View September 27, 2020 15:11
Start today EVS B-Ionic Nitrate add 3ml/day for 5 days. Test No3 on fifth day. Current No3 is 2ppm
View September 16, 2020 20:40
x2 ORA Purple plasma died, unknown reason why atm....

Lightsaber Milli is still okay but seems less colorful.

Noticed ATO Sensor was not working so I cleaned off and it worked. Need to send it back to manufacture.
-------> Salinty 1.032 Calibration Solution wrong for Refractometer.
View August 7, 2020 04:19
Stored 1.026 Sg water mixed 13 days ago tested 1.026 Sg and 12.82 dkh.
Added 5 Frags from CVR member Matt Fugman. (x2)ORA Purple Plasma, Forest Fire Digi, Lightsaber Milli, Tableing acro.
.
PLAN: Watch Corals for 2 weeks and if it looks like they are not doing good then add more flow.
View August 4, 2020 03:35